# Entropy Heat-Mapping: Localizing GPT-Based OCR Errors with Sliding-Window Shannon Analysis

### v1

@@ Abstract
Vision–language models such as OpenAI GPT-4o can transcribe mathematical documents directly from images, yet their token-level confidence signals are seldom used to pinpoint local recognition mistakes. We present an entropy-heat-mapping proof-of-concept that turns per-token Shannon entropy into a visual “uncertainty landscape.” By scanning the entropy sequence with a fixed-length sliding window, we obtain hotspots that are likely to contain OCR errors such as missing symbols, mismatched braces, or garbled prose. Using a small, curated set of scanned research pages rendered at several resolutions, we compare the highlighted hotspots with the actual transcription errors produced by GPT-4o. Our analysis shows that the vast majority of true errors are indeed concentrated inside the high-entropy regions. This study demonstrates—in a minimally engineered setting—that sliding-window entropy can serve as a practical, lightweight aid for post-editing GPT-based OCR. All code, sample data, and annotation guidelines are released to encourage replication and further research.
1 Background and Related Work
This section provides context on (i) the information–theoretic lens for understanding OCR, (ii) the extension of GPT from purely text-based to vision–language modeling, (iii) prior efforts leveraging per-token probability distributions for confidence checks, and (iv) existing methods for local error detection in related domains. We then outline where our sliding-window entropy framework fits within this broader literature, especially for mathematical OCR tasks.
@@ 1.1 Information Theory and OCR
Optical Character Recognition (OCR) can be analyzed through the prism of a noisy channel model. In such a view, an image $X$ (carrying text or mathematical symbols) is transmitted through a channel that might introduce noise via image blur, compression artifacts, or similar distortions. The output of that channel, a recognized text sequence $Y$, reflects an imperfect reconstruction of the original message. Foundational concepts[1] like entropy capture the overall uncertainty or “information content” in $X$ and $Y$. With higher resolution or lighter noise, $Y$ is more faithful to $X$ (i.e., $I(X;Y)$ is larger, and $H(Y\mid X)$ is smaller), implying fewer OCR errors.
Prior work has explored uncertainty quantification for text recognition using modern techniques like beam-search and dropout ensembles. For instance, Patel et al. [2] introduced Seq-UPS, a semi-supervised framework that leverages dropout-based uncertainty estimates to select reliable pseudo-labels, yielding improved OCR performance. Fang et al. [3] likewise proposed a Dropout Decoding method for large vision–language models; by explicitly discarding the most uncertain visual tokens, they reduced error rates on multimodal tasks such as visual question answering. Mathematical expressions, in particular, can be sensitive to OCR distortions: a single misread Greek symbol or missing bracket can disrupt an entire formula. Localizing such errors requires more fine-grained uncertainty signals, which motivates the present research.
@@ 1.2 GPT for Vision–Language Tasks
Generative pre-trained transformers (GPT) have evolved to handle multimodal inputs, merging deep text representations with image embeddings [4, 5]. Instead of restricting transformers to textual corpora, these extended architectures accept both an image $X$ and a textual prompt and generate an output token stream, $y_{1},y_{2},\dots,y_{n}$, from a vocabulary $\mathcal{V}$. Specifically,
$P(y_{i}\mid X,y_{1},\ldots,y_{i-1}),$
encodes the conditional probability of each new token given the image and preceding output tokens.
One promising application of these multimodal GPT models is mathematical OCR, where LaTeX transcription requires fine visual parsing. Luo and Xu [6] addressed contextual bias in transformer-based OCR decoders, emphasizing the importance of visual inputs for ambiguous symbols. Chen et al. [7] proposed a Recognition and Language Fusion Network that uses a lightweight language component alongside a vision core, significantly boosting math transcription accuracy on CROHME benchmarks. Similarly, Hyeon et al. [8] combined audio-based recognition with a small LM to rival GPT-4’s math OCR accuracy for lecture transcriptions. Going beyond purely mathematical text, Kim et al. [9] showed that large multimodal LLMs (e.g. GPT-4 Vision) can sometimes outperform classical OCR engines on difficult handwriting tasks, though integrative approaches that fuse external OCR signals can yield additional gains.
1.3 Token–Level Probability and Entropy
Modern large language model APIs commonly return a truncated set of top-$k$ tokens with associated log-probabilities at each time step. Converting this partial distribution into Shannon entropy (plus a single “tail” bucket for all lower-ranked tokens) offers a practical lower-bound estimate of the model’s true uncertainty at each position. This truncated entropy often suffices to detect peaks of uncertainty, especially if the tail probability is relatively small.
Recent advances in black-box LLM calibration suggest that model-generated confidences are themselves prone to miscalibration. Xiong et al. [10] found that naive confidence prompts in closed-source LLMs can be highly overconfident, while Yang et al. [11] proposed using multiple rephrasings of a query to improve answer consistency and estimate model uncertainty. Tonolini et al. [12] introduced Bayesian Prompt Ensembles to blend multiple prompts, producing more stable uncertainty measures with lower expected error. Yona et al. [13] further highlight that even aligned LLMs often sound more certain than their internal distribution suggests, while Wang et al. [14] provide a curated benchmark (UBench) to systematically test these models’ calibration on multiple-choice tasks.
Likewise for OCR, confidence gleaned from token-level distribution can guide error detection and correction. Although OCR often aggregated such signals over entire sequences, the ability to track local spikes in entropy opens clearer paths to targeted proofreading, as a single confusion can prove fatal for some mathematical expressions.
1.4 Localising Errors via Sliding Windows
In related fields like neural machine translation and speech transcription, researchers have experimented with short segment-level confidence measures to detect local mistakes. They often slice text into small windows and flag segments where the model’s probability mass is notably scattered. While this concept is relatively unexplored for GPT-based mathematical OCR, some studies demonstrate that local error-detection can substantially improve text quality across languages and scripts. For instance, Rigaud et al. [15] organized a post-OCR correction that split detection and correction into separate tasks, and Chen and Ströbel [16] proposed a hybrid method for post-processing outputs with a character-aware BERT (Bidirectional Encoder Representations from Transformers), pinpointing localized errors in historical handwriting. Iterative refinement [17] and two-stage pipelines [18] likewise demonstrate that targeting suspicious segments can outperform naive single-pass systems. Soper et al. [19] and Boros et al. [20] both show how large language models can be adapted for post-OCR text correction—though they also caution that out-of-the-box LLMs, while powerful, are not always reliable fixers of OCR errors. Bourne [21] made strides by synthesizing artificial errors to fine-tune a language model, cutting error rates even more effectively than some real-data approaches.
Within mathematical domains, local segmentation is especially relevant: a single bracket or minus sign can be misread, yet compromise the entire expression’s correctness. Our work introduces an entropy heat-mapping pipeline that computes a sliding-window average of per-token entropy. Regions with high mean entropy become hotspots for review—mirroring localized error detection in machine translation or historical text correction, but now tailored to GPT-based math OCR.
@@ 2.5 Key Model Assumptions
    Top-$k$ Access. APIs provide only the highest-probability tokens; we approximate $H_{i}$ with the truncated‐entropy lower bound, sufficiently accurate when the tail mass is small.

### v2

@@ Abstract
Vision-language models such as OpenAI GPT-4o can transcribe mathematical documents directly from images, yet their token-level confidence signals are seldom used to pinpoint local recognition mistakes. We present an entropy-heat-mapping proof-of-concept that turns per-token Shannon entropy into a visual ”uncertainty landscape”. By scanning the entropy sequence with a fixed-length sliding window, we obtain hotspots that are likely to contain OCR errors such as missing symbols, mismatched braces, or garbled prose. Using a small, curated set of scanned research pages rendered at several resolutions, we compare the highlighted hotspots with the actual transcription errors produced by GPT-4o. Our analysis shows that the vast majority of true errors are indeed concentrated inside the high-entropy regions. This study demonstrates–in a minimally engineered setting–that sliding-window entropy can serve as a practical, lightweight aid for post-editing GPT-based OCR. All code and annotation guidelines are released to encourage replication and further research.
Keywords: Sliding-window entropy, Mathematical OCR, Token-level uncertainty heat-mapping, Error localisation.
1 Introduction and Related Work
Transcribing mathematical documents from scanned images poses unique challenges, as subtle symbol confusion (e.g., minus signs vs. tildes) can undermine entire expressions. In this section, we review three broad research directions: (i) information–theoretic perspectives on OCR, (ii) the emergence of GPT for multimodal tasks, and (iii) prior efforts at uncertainty-based error detection—both global and local. We then show how our sliding-window entropy framework extends these lines of work to pinpoint where within a GPT-based transcript recognition mistakes arise, with a particular focus on mathematical OCR.
@@ 1.1 Information Theory and OCR
Optical Character Recognition (OCR) can be framed through the lens of a noisy-channel model [1], where an image $X$ containing text must be decoded despite potential blur, compression artifacts, or other noise sources—resulting in an imperfect output sequence $Y$. Classical information measures (e.g., entropies $H(X)$ and $H(Y)$, or the mutual information $I(X;Y)$) help describe how uncertainty in $X$ propagates to $Y$. Modern research has further explored ways to quantify OCR uncertainty and mitigate errors. For instance, Patel et al. [2] derive per-sequence uncertainty estimates via Monte Carlo dropout to filter unreliable pseudo-labels in text recognition. Fang et al. [3] describe a dropout-decoding strategy in large vision–language models, illustrating that high global uncertainty frequently correlates with poor recognition accuracy. These approaches demonstrate that confidence measures can be leveraged to improve OCR outputs—particularly for complex tasks like mathematical symbol detection, where even a single misread bracket may invalidate an entire formula.
@@ 1.2 GPT for Vision–Language Tasks
Over the past few years, generative pre-trained transformers have evolved from purely text-based systems into multimodal networks that interpret images alongside textual prompts [4, 5]. These extended GPT engines map image embeddings and partially decoded text into probabilities over a large vocabulary. Vision–language GPT models have proven especially promising for mathematical OCR, where long-range context is needed to disambiguate complex notation. Luo and Xu [6], for example, emphasize the role of precise visual cues to differentiate visually similar Greek letters, while Chen et al. [7] integrate a lightweight language model with a dedicated visual core to reduce transcription errors on CROHME math benchmarks. Recent studies also show that GPT’s visual modules can handle challenging inputs like handwritten math [8] or audio-based lecture notes [9], sometimes outperforming traditional OCR pipelines. Nevertheless, even top-tier vision–language models can generate localized mistakes—missing or confusing symbols—in ways that standard system-level metrics fail to capture.
1.3 Token-Level Entropy versus Global Uncertainty
When large language models (LLMs) generate text, their APIs often provide the top-$k$ tokens with corresponding log-probabilities at each decoding step. Summing across these probabilities yields a truncated distribution suitable for approximate Shannon entropy. Several works have noted that peaks in token-level entropy can reflect points of confusion in text generation [10, 11, 12, 13, 14]. For OCR, such confusion likely stems from visually ambiguous inputs: a blurred superscript, a partially occluded minus sign, and so on.
Yet, many existing systems focus on global or sequence-level confidence. For instance, Wang et al. [15] calibrate a single “energy-based” score for visual classifiers that flags uncertain images, while Chun et al. [16, 17] embed entire image–text pairs into probabilistic spaces and measure differential entropy as a reliability indicator. Similarly, Upadhyay et al. [18, 19] attach Bayesian adapters to frozen networks, using overall negative log-evidence or variance as a global confidence metric to detect untrustworthy predictions. Fei et al. [20] rank candidate captions by their total sequence-level entropy to identify robust image captions, and Ji et al. [21] monitor global KL-divergence to decide when multimodal embeddings are stable. For cross-modal retrieval tasks, Li et al. [22] reduce all Dirichlet-based evidence into a single score, and Khan & Fu [23] show that the overall distributional entropy correlates with unreliable VQA results. Zhou et al. [24] also rely on total entropy to spot ambiguous event-based actions. While these global measures can signal that an entire output is potentially flawed, they provide no direct insight on which part of the text is at fault. By contrast, we exploit local per-token signals to automatically pinpoint suspicious sub-sequences—crucial for post-editing mathematical text, where minor symbol mismatches can cause outsized damage.
1.4 Local Error Detection via Sliding Windows
To address the problem of where errors occur, researchers in machine translation or speech recognition often slice transcripts into short windows and measure local confidence. Rigaud et al. [25] propose a post-OCR correction process that operates on small segments, while Chen and Ströbel [26] refine OCR output by selectively re-checking suspicious tokens with a character-aware language model. Similar segment-wise or iterative approaches have been adopted for historical text [27, 28], and indeed large language models like BART have been retrained to patch local OCR mistakes [29, 30]. Bourne [31] goes further by synthesizing artificial errors in order to robustly fine-tune correction models, reducing error rates more effectively. Notably, while these pipelines vary, they share a common theme: error-prone sub-sequences can be more effectively targeted—either for human proofing or re-prompted corrections—than scanning an entire document blindly.
In the context of mathematical OCR, local segmentation is especially critical: dropping a subscript underscore or mixing up braces can invalidate a formula. Our work introduces entropy heat-mapping as a lightweight technique to identify hotspots of local uncertainty. Specifically, we (i) compute truncated entropy for each token, (ii) apply a sliding window to produce interval-based averages, and (iii) highlight the windows with highest mean entropy. These windows correlate strongly with actual symbol confusion or missing tokens, as demonstrated in our proof-of-concept results.
1.5 Our Contribution
Although several recent studies [15, 16, 18, 20, 21, 22, 23, 24] employ global entropy or other confidence scores to gauge the overall reliability of vision–language models, the question of where recognition errors arise inside a generated transcript remains largely unexplored. The present paper addresses that gap by proposing an entropy-based localisation framework for GPT-driven mathematical OCR.
Our main contributions are:
  1. 1.
    Sliding-window Shannon analysis.  We transform the per-token log-probabilities returned by modern GPT APIs into a linear-time, sliding-window entropy signal that yields a fine-grained “uncertainty landscape” over the output text.
  2. 2.
    Entropy heat-mapping for error detection.  Visualising the windowed entropies highlights short contiguous spans—hotspots—whose elevated uncertainty correlates strongly with actual transcription mistakes such as missing symbols, mismatched braces, or dropped subscripts.
  3. 3.
    Instantiating the classical coarse-graining property for top-$k$+tail entropy.  It has long been known in information theory (e.g. Cover & Thomas, Sec. 2) that coarse-graining—merging several outcome categories into one—can only decrease Shannon entropy. We adapt this classical fact to the practical GPT setting in which APIs expose probabilities only for the top-$k$ tokens, collapsing the remaining mass into a single “tail” bucket. Our short, self-contained derivation shows exactly how the coarse-graining inequality applies to per-token log-probabilities and clarifies its implications for black-box uncertainty extraction in vision–language OCR.
  4. 4.
    Proof-of-concept human evaluation.  On a curated mini-corpus of scanned research pages rendered at three resolutions (72, 150, and 300 dpi), we demonstrate that the vast majority of human-verified OCR errors fall inside the few highest-entropy windows, validating the practical utility of the proposed heat-mapping approach.
  5. 5.
    Open-source resources.  All code, sample images, entropy-extraction scripts, and detailed annotation guidelines are publicly released to encourage replication and further research on local uncertainty cues in multimodal GPT systems.
In essence, the paper converts raw probabilistic signals—already available from mainstream GPT endpoints—into an actionable tool for post-editing: it pinpoints precisely where the model is least certain and therefore where human reviewers or automatic re-prompting should focus.
Reproducibility.
The complete implementation and annotation protocol are freely available at https://github.com/Alexei-WLU/scan2latex-entropy.
1.6 Organization of the Paper
We formally define the local uncertainty detection problem in Section 2, specifying how GPT-based OCR outputs and per-token probabilities lead to window-level entropies. Next, Section 3 presents our sliding-window technique, including exact computation of top-$k$+tail entropy and approaches for highlighting hotspots in linear time. In Section 4, we detail the proof-of-concept experiments and human annotation protocol, while Section 5 reports qualitative and quantitative results. We discuss broader implications and limitations in Section 6, and Section 7 concludes with future prospects for scaling entropy heat-mapping to real-world OCR scenarios.
@@ 2.5 Key Model Assumptions
    Top-$k$ Access. APIs provide only the highest-probability tokens; we approximate $H_{i}$ with the truncated-entropy lower bound, sufficiently accurate when the tail mass is small.
@@ 6.4 Broader Implications
7 Conclusion and Future Work
In this paper, we presented a novel proof-of-concept framework for localizing OCR errors in GPT-based mathematical transcripts using per-token entropy signals. By converting per-token log-probabilities into a practical “uncertainty landscape,” our method pinpointed short contiguous segments—hotspots—most likely to contain recognition problems. The sliding-window averaging approach proved to be a lightweight and interpretable means of detecting local failures, as demonstrated on a small-scale corpus of scanned research pages rendered at several resolutions.
Qualitative evaluation showed that high-entropy windows align with various OCR pitfalls, including missing symbols, mismatched delimiters, and confusions among visually similar characters. Although our findings relied on human annotation over a modest set of 36 image samples, the strong anecdotal evidence suggests that this uncertainty-driven technique can help researchers or proofreaders focus attention on precisely those sub-sequences that merit closer review. Notably, it operates with minimal overhead, given that most GPT-based APIs already expose top-k token probabilities.
Nonetheless, we acknowledge certain limitations. First, the overall study size remains small, so additional experiments with larger and more diverse datasets are necessary to confirm the generality of our insights. Second, our “hotspot” detection occasionally flagged high-entropy regions that did not contain outright errors, pointing to a need for improved thresholds or follow-up diagnostics (e.g., automated re-prompting). Finally, the method was tested primarily on text with stationary mathematical expressions—tables, complex layouts, or heavily multi-column documents might require more advanced segmentation and tracking techniques.
Looking ahead, several extensions and directions emerge:
  • Automated Re-Prompting. The proposed hotspot detection can be combined with a selective re-querying of GPT, requesting localized corrections only for suspicious subsequences.
  • Fine-Grained Error Classification. Beyond marking high-uncertainty tokens as potential errors, a downstream classifier could distinguish among error types (e.g., missing symbol vs. symbol confusion) to inform specialized correction modules.
  • Adaptive Windowing. Varying the window length automatically in response to local text features (such as formula boundaries) could enhance precision in highlighting exactly the relevant tokens.
  • Multi-Task and Multilingual Extensions. The same sliding-window entropy approach may prove useful in other scenarios (e.g., handwriting recognition, multilingual OCR) where local token competition signals confusion.
  • Larger-Scale Benchmarking. A more expansive and systematic data collection—covering more diverse page layouts, fonts, and scanning artifacts—would facilitate rigorous quantitative comparisons with alternative error-spotting methods.
In conclusion, our entropy heat-mapping framework illustrates how local confidence signals can offer a powerful lens into GPT-based OCR pipelines. By exposing precisely which tokens are most uncertain, it sets the stage for deeper refinements and targeted corrections, thereby improving the reliability of OCR outputs while minimizing human proofreading workloads. We hope that this proof-of-concept fosters further research on token-level uncertainty in large language models and drives the development of more reliable, transparent multimodal transcription systems.
Declarations
Conflict of Interest
The authors have no relevant financial or non-financial interests to disclose. The authors have no conflict of interest to declare for the content of this article.
Code Availability
The Python code used in this study, developed by the authors, is openly accessible on GitHub at https://github.com/Alexei-WLU/scan2latex-entropy. This repository contains all scripts necessary to reproduce the experimental results presented in this paper.